OpenStudy (anonymous):

5 + 2 = 7 Choose the related addition sentence that shows the commutative property of addition.

OpenStudy (anonymous):

5 + 2 = 2 + 5

OpenStudy (anonymous):

8 + 7 = 15 Choose the related addition sentence that shows the commutative property of addition.

OpenStudy (anonymous):

8 + 7 = 7 + 8.., it easy.., just flip first number to second, and second to first